### v2

# Open ASR Leaderboard: Towards Reproducible and Transparent Multilingual and Long-Form Speech Recognition Evaluation

Abstract
Despite rapid progress, ASR evaluation remains saturated with short-form English, and efficiency is rarely reported. We present the Open ASR Leaderboard, a fully reproducible benchmark and interactive leaderboard comparing 60+ open-source and proprietary systems across 11 datasets, including dedicated multilingual and long-form tracks. We standardize text normalization and report both word error rate (WER) and inverse real-time factor (RTFx), enabling fair accuracy–efficiency comparisons. For English transcription, Conformer encoders paired with LLM decoders achieve the best average WER but are slower, while CTC and TDT decoders deliver much better RTFx, making them attractive for long-form and offline use. Whisper-derived encoders fine-tuned for English improve accuracy but often trade off multilingual coverage. All code and dataset loaders are open-sourced to support transparent, extensible evaluation.
Index Terms—  Benchmarking, automatic speech recognition, reproducible, multilingual, long-form
1 Introduction
The expression “jack of all trades, master of none” is typically used to point out someone who has not gained expertise in a specific task. Interestingly, the second part of the original phrase – “though oftentimes better than master of one” – is usually omitted, which altogether changes its meaning. This tradeoff is highly relevant for machine learning systems: should a model excel at a single task, or perform competently across multiple tasks?
Automatic speech recognition (ASR) has seen remarkable progress in recent years, fueled in part by open-source contributions. Publicly-available datasets [1, 2] and pre-trained models [3, 4, 5] have enabled researchers across academia and industry to build on existing work. Yet, as the number of datasets and models grows, it becomes increasingly difficult for developers of new models to know which baselines to compare against and how. Similarly, users focused on inference may find it difficult to identify which model, whether open-source or proprietary, best meets their needs in terms of application and/or efficiency. Moreover, most existing benchmarks and evaluations overwhelmingly emphasize English and short-form transcription.
Several efforts have sought to address parts of this problem, including benchmarks across multiple accents and diverse contexts in the French language [6], under noise and reverberation in far-field settings [7], comparing commercial and open-source models in English and Chinese [8]. Some common observations can be drawn from these efforts: (1) there is no “catch-all” model, (2) no single dataset is sufficient for evaluation, and (3) a single metric, i.e., word error rate (WER), is not enough.
To address these challenges we introduce the Open ASR Leaderboard. Our contributions include:
  1. 1.
    An interactive leaderboard that compares $60+$ open-source and proprietary models from $18$ organizations, with evaluations over $11$ datasets. [^1] Leaderboard: hf.co/spaces/hf-audio/open_asr_leaderboard
  2. 2.
    A multilingual benchmark covering German, French, Italian, Spanish, and Portuguese.
  3. 3.
    A comparison of models for long-form transcription.
For full transparency and to facilitate the addition of new models and datasets, the leaderboard’s codebase is open-sourced. [^2] Code: github.com/huggingface/open_asr_leaderboard The above model, dataset, and languages count are as of 8 Oct 2025, and will continue to increase with new additions to the leaderboard.
2 Open ASR Leaderboard
Table 1: Datasets used for the Open ASR Leaderboard. The Task(s) column indicates which tasks (as denoted in Section 2.1) the dataset is used for. The duration is of the test-split. The Multilingual datasets indicate the range of durations for the evaluated languages. Note that for Earnings22 a subset of $5\text{\,}\mathrm{h}$ is used for short-form English (Leaderboard) comparison.
Dataset Task(s) Duration [h] License Source Style Transcriptions
AMI Meeting Corpus [9] Leaderboard 9 $C$-BY-4.0 Meetings Spontaneous Punctuated, cased
CoVoST-2 [10] Multilingual (de/fr/it/es/pt) 5.3–23 $C$-BY-NC-4.0 Open domain Read Punctuated, cased
Earnings21 [11]
Long-form
39 $C$-BY-SA-4.0 Earnings calls Oratory, spontaneous Punctuated, cased
Earnings22 [12]
Leaderboard, Long-form
119 $C$-BY-SA-4.0 Earnings calls Oratory, spontaneous Punctuated, cased
FLEURS [13] Multilingual (de/fr/it/es/pt) 2.0–3.5 $C$-BY-4.0 Wikipedia Read Punctuated, cased
GigaSpeech [2] Leaderboard 40 apache-2.0 Audiobook, podcast, YouTube Read, spontaneous Punctuated
LibriSpeech (clean) [1] Leaderboard 5.4 $C$-BY-4.0 Audiobooks Read Normalized
LibriSpeech (other) [1] Leaderboard 5.1 $C$-BY-4.0 Audiobooks (noisier) Read Normalized
MLS [14] Multilingual (fr/it/es/pt) 0.8–6.3 $C$-BY-4.0 Audiobooks Read Normalized
SPGISpeech [15] Leaderboard 100 User Agreement Financial meetings Oratory, spontaneous Punctuated, cased
TED-LIUM v3 [16]
Leaderboard, Long-form
3 CC-BY-NC-ND 3.0 TED Talks Oratory Normalized
VoxPopuli [17] Leaderboard 5 CC0 European Parliament Oratory Punctuated
2.1 Overview
As part of the Open ASR Leaderboard, there are evaluations on three tasks, with the following tabs:
  1. 1.
    Leaderboard, which evaluates English transcription.
  2. 2.
    Multilingual, which currently evaluates German, French, Italian, Spanish, and Portuguese transcription.
  3. 3.
    Long-form, which evaluates English transcription on audio longer than $30\text{\,}\mathrm{s}$ (i.e., the receptive field of Whisper [3]).
A separate Long-form evaluation is necessary because some models employ chunking strategies to reduce inference time, which can in turn affect transcription quality.
The datasets used for evaluation are presented in Section 2.2. The models are evaluated according to WER, and can be dynamically sorted on the leaderboard page according to WER-performance on a particular dataset (more on metrics in Section 2.3). Section 2.4 presents models that are evaluated within the Open ASR Leaderboard, while Section 2.5 describes the process of adding a new model.
2.2 Datasets
Table 1 summarizes the datasets used for the Open ASR Leaderboard.
Dataset retrieval and usage is enabled through the datasets package [18]. The datasets themselves are hosted on the Hugging Face Hub, [^3] Leaderboard: hf.co/datasets/hf-audio/esb-datasets-test-only-sorted;
Multilingual: hf.co/datasets/nithinraok/asr-leaderboard-datasets;
Long-form: hf.co/datasets/hf-audio/asr-leaderboard-longform
which allows for interactive exploration, e.g., listening to individual audio, viewing their metadata, and performing SQL queries, all from the browser and without having to download the datasets. With the above engineering choices, the datasets can be conveniently downloaded and used in Python, as shown below:
Listing 1: Example of loading dataset.
from datasets import load_dataset
ds = load_dataset("hf-audio/esb-datasets-test-only-sorted", "ami", split="test")
audio_sample = ds[0]
2.3 Metrics
We report results on two metrics: word error rate (WER) for comparing transcription quality, and inverse real-time factor (RTFx) for comparing inference speed.
To account for differences in punctuation and casing between model outputs and the dataset transcriptions (see Table 1), we normalize text before computing WER. Punctuation and casing are removed, and an English text normalization pipeline, closely following Whisper [3], is applied. This includes number normalization (e.g., “0” to “zero”), spelling standardization, and filler word removal (e.g., “uh”, “mhm”).
RTFx is defined as:
$\text{RTFx}=\frac{\text{Total duration of audio}}{\text{Transcription time}},$ (1)
with a higher value being better (i.e., lower latency). RTFx can be computed on a single audio or a batch. RTFx (and variants) is a useful metric for quantifying a model’s efficiency in processing long audio samples [19].
2.4 Current models
Table 2: Distribution of encoder and decoder architectures for the open-source models in the Open ASR Leaderboard. Some models use hybrid architectures for the encoder or decoder, and are counted twice.
Enc $\downarrow$ / Dec $\rightarrow$ Transformer CTC RNN-T/TDT LLM Total
Conformer-based 5 7 8 4 24
Whisper 18 0 0 2 20
Self-supervised 0 13 0 0 13
Custom 3 0 0 0 3
Total 26 20 8 6 60
Table 3: Subset of Open ASR Leaderboard results on English transcription. WER is averaged over datasets corresponding to the Leaderboard in Table 1. Whisper-FT stands for Whisper-finetuned.
Model Avg. WER $\downarrow$ Avg. WER rank RTFx $\uparrow$ Open Encoder Decoder # Lang.
NVIDIA Canary Qwen 2.5B 5.63 1 418.28 Yes FastConformer [20] LLM-based 1
IBM Granite Speech 3.3 8B 5.74 2 145.42 Yes Conformer [21] LLM-based 5
IBM Granite Speech 3.3 2B 6.00 3 259.57 Yes Conformer [21] LLM-based 5
Microsoft Phi 4 Multimodal Instruct 6.02 4 151.1 Yes Conformer [21] LLM-based 8
NVIDIA Parakeet TDT 0.6B v2 6.05 5 3386.02 Yes FastConformer [20] TDT [22] 1
Aqua Voice Avalon 6.24 6 - No - - 1
NVIDIA Parakeet TDT 0.6B v3 6.32 7 3332.74 Yes FastConformer [20] TDT [22] 25
NVIDIA Canary 1B Flash 6.35 8 1045.75 Yes FastConformer [20] Transformer 4
Kyutai STT 2.6B en 6.40 9 88.37 Yes Mimi codec [23] Transformer 1
NVIDIA Canary 1B 6.50 10 235.34 Yes FastConformer [20] Transformer 4
Nyra Health CrisperWhisper 6.67 11 84.05 Yes Whisper-FT [24] Whisper-FT 1
ElevenLabs Scribe v1 6.88 12 - No - - 99
Speechmatic Enhanced 6.91 13 - No - - 55
Mistral AI Voxtral Mini 3B 7.05 16 109.86 Yes Whisper-FT [25] LLM-based 8
RevAI Fusion 7.12 18 - No - - 1
NVIDIA Canary 1B v2 7.15 20 749 Yes FastConformer [20] Transformer 25
Distil-Whisper Large v3.5 7.21 21 202.03 Yes Whisper [3] Transformer 1
NVIDIA Parakeet CTC 1.1B 7.40 23 2728.52 Yes FastConformer [20] CTC 1
OpenAI Whisper Large v3 7.44 25 145.51 Yes Whisper [3] Whisper 99
NVIDIA FastConformer CTC Large 8.96 44 6399.25 Yes FastConformer [20] CTC 1
SpeechBrain ASR wav2vec2 Librispeech 14.4 52 451.18 Yes wav2vec2 [4] CTC 1
Of the 64 models currently listed in the Open ASR Leaderboard (as of 8 Oct 2025), 57 are open-source. The 64 models come from 18 organizations: NVIDIA (18), Meta (12), OpenAI (8), Hugging Face (5), University of Washington (4), IBM (2), Rev AI (2), SpeechBrain (2), Useful Sensors (2), AssemblyAI (1), Aqua Voice (1), ElevenLabs (1), Kyutai (1), Microsoft (1), Mistral AI (1), Nyra Health (1), Speechmatics (1), and Ultravox (1).
Table 4: Average WERs for different languages (German/French/Italian/Spanish/Portuguese) on the Multilingual datasets in Table 1.
Model DE FR IT ES PT
Microsoft Phi 4 Multimodal Instruct 4.50 5.13 4.80 3.59 5.15
NVIDIA Canary 1B v2 4.96 4.86 5.66 3.22 6.23
OpenAI Whisper Large v3 4.97 6.59 5.14 3.32 4.38
NVIDIA Parakeet TDT 0.6B v3 4.90 5.38 5.58 3.72 5.95
Mistral AI Voxtral Mini 3B 5.36 5.96 5.88 3.81 4.80
ElevenLabs Scribe v1 10.8 15.1 7.71 9.28 22.8
Table 2 provides a breakdown of the encoder and decoder architectures of the models in the leaderboard, while some of the models are presented in our results in Tables 3, 4 and 5. Conformer-based encoders [21, 20], Whisper-based encoders [3], self-supervised encoders (i.e., wav2vec2 [4], HuBERT [5], data2vec [26]) and custom encoders [23, 27] are considered. Whisper-based encoders either use the encoder model as is [28], apply low-rank adaptation [29], or fine-tune [24, 25]. As decoders, the following are represented: transformed-based, CTC (Connectionist Temporal Classification), Recurrent Neural Network Transducer (RNN-T), Token-and-Duration Transducer (TDT) [22], and LLM (Large Language Model)-based. While LLM-based decoders are transformer-based, we denote a separate category for approaches that use a pre-trained LLM.
2.5 Adding a new model or dataset
All the evaluation code is open-sourced on GitHub.444github.com/huggingface/open_asr_leaderboard As such, the process of adding a new model or dataset consists of opening a pull request (PR). To add a new model, a new folder should be created with: (1) a Python script for evaluating on a specific dataset and (optionally) with a specific model version; [^5] A template can be found in transformers/run_eval.py and (2) a Bash script which calls the Python script for each dataset and model combination.
3 Results
All evaluation scripts, as described in Section 2.5, were conducted on an NVIDIA A100-SXM4-80GB GPU (driver 560.28.03, CUDA 12.6), using a batch size of 64 whenever memory allowed, and reduced adaptively (48, 32, 16, …) when necessary to fit in device memory. Since the full results are continuously updated on the Open ASR Leaderboard and are too extensive to include here, we present a condensed version of the English leaderboard in Table 3, the multilingual results in Table 4, and the long-form results in Table 5.
For short-form English transcription (Table 3), models with a Conformer encoder and an LLM-based decoder achieve the best results, with the top four models using this architecture. However, these approaches are significantly slower compared to models using TDT or CTC decoders. While the latter can achieve superior RTFx, this comes at the cost of accuracy: e.g., the best CTC-based model (NVIDIA Parakeet CTC 1.1B) ranks only 23rd in terms of WER.
Self-supervised learning (SSL) approaches have enabled ASR systems for 1K+ languages, yet the top SSL-based system for English transcription ranks only 52nd, i.e., last row of Table 3 which uses wav2vec2 [4]. As shown in Table 2, the leaderboard only covers SSL encoders with CTC decoders; combining such encoders with more performant decoders may help bridge this gap.
Another popular trend is to leverage Whisper’s encoder (see Table 2), which has been trained on a large multilingual corpus. As shown in Table 3, models that fine-tune Whisper’s encoder [24, 25] (or train a new decoder [28]) achieve better average WER than OpenAI Whisper Large v3.
The multilingual results of the Open ASR Leaderboard (Table 4) highlight an important tradeoff in ASR: specialization versus broad coverage. Namely, the improvement in English performance of fine-tuned Whisper models often comes at the cost of multilingual coverage: Whisper-derived models typically train on fewer languages (often just English), while Whisper itself supports 99 languages. Similarly, NVIDIA’s models demonstrate this tradeoff more clearly: Parakeet TDT 0.6B v3 adds multilingual support compared to v2, and Canary 1B v2 expands from 4 to 25 languages. In both cases, broader language coverage comes at the cost of English transcription accuracy.
For Tables 3 and 4 (short-form and multilingual), open-source models show the strongest performance. The highest-ranking closed-source model (Aqua Voice Avalon) ranks 6th. Fairly computing RTFx for closed-source models is not possible due to upload latency and lack of GPU usage control, making direct efficiency comparisons infeasible.
Table 5: Average WER (%) and RTFx on the Long-form datasets of Table 1.
Model Avg. WER Avg. WER rank RTFx
ElevenLabs Scribe v1 4.33 1 -
RevAI Fusion 5.04 2 -
Speechmatics Enhanced 5.08 3
OpenAI Whisper Large v3 6.43 6 68.56
Kyutai STT 2.6B en 6.64 8 52.56
NVIDIA Parakeet CTC 1.1B 6.68 9 2793.75
NVIDIA Parakeet TDT 0.6B v2 6.91 10 955.87
Distil-Whisper Large v3 7.27 14 153.26
NVIDIA FastConformer CTC Large 11.14 17 5531.13
On the other hand, for long-form transcription (Table 5), closed-source systems, such as ElevenLabs and RevAI, achieve the lowest error rates, likely benefiting from domain-specific tuning and production-grade infrastructure. Among open-source models, OpenAI Whisper Large v3 is the strongest, with distilled variants providing faster inference. As in the short-form setting, NVIDIA’s CTC- and TDT-based models significantly improve throughput with only moderate losses in quality, making them particularly well suited for large-scale or offline transcription of lengthy audio.
4 Conclusion
We present the Open ASR Leaderboard, a reproducible benchmark covering $60+$ systems and $11$ datasets, including multilingual and long-form speech. Standardized text normalization enables a unified basis for comparing WER performance accuracy, and our RTFx evaluation allows for efficiency comparisons. Conformer–LLM models achieve the strongest English WER but at the cost of higher latency, whereas CTC/TDT decoders offer faster inference with only modest accuracy trade-offs, making them attractive for long-form transcription. All code and datasets are fully open-sourced to support transparent and extensible evaluation.
Future work include expanding language and domain evaluations (e.g., far-field speech), incorporating additional metrics (e.g., token error rate [8]), and exploring underrepresented encoder–decoder combinations (see Table 2). With the rise of LLMs and their ability for strong ASR performance (Table 3), we expect more approaches that employ them.
ACKNOWLEDGMENT
The authors would like to thank all the contributors to the Open ASR Leaderboard.